OpenStudy (anonymous):

What is the solution to the equation log 3 x – log 3 2 = 2 ? 3 is the base.

OpenStudy (anonymous):

log3(x/2)=2 continue from here

OpenStudy (shubhamsrg):

with equal base,, log x + log y = log (xy)

OpenStudy (shubhamsrg):

it should be easier now..also log (base b) a = c => b^c = a this should help